OpenStudy (anonymous):

Andy estimated that he would need 72 feet of lumber for a tree house project. He later found that the actual amount of lumber needed was 61 feet. What was the percent error of Andy's estimation?

OpenStudy (texaschic101):

72 - 61 = 11 11/61 = 0.18 = 18%

OpenStudy (anonymous):

ty

OpenStudy (texaschic101):

anytime :)

OpenStudy (anonymous):

A CD usually sells for $15.00. If the CD is 10% off, and sales tax is 7%, what is the total price of the CD, including tax?

OpenStudy (anonymous):

@texaschic101

OpenStudy (texaschic101):

15 - .10(15) = 15 - 1.50 = 13.50 13.50 (.07) = 0.945 rounds to 0.95 total price is : 14.45

OpenStudy (anonymous):

ok last one A pair of shoes usually sells for $69. If the shoes are 40% off, and sales tax is 7%, what is the total price of the shoes, including tax?

OpenStudy (texaschic101):

69 - .40(69) = 69 - 27.60 = 41.40 41.40 (.07) = 2.898 rounds to 2.90 total cost is : 41.40 + 2.90 = 44.30